# Avoiding Catastrophe in Online Learning by Asking for Help

### v3

@@ B.2 Main proof
Next, fix an $a\in A$ and let $S_{a}=\{s\in\mathbf{s}:\pi^{m}(s)=a\}$ be the set of observed states which share a mentor action. We claim that $\hat{S}_{2}\cap S_{a}$ is a packing of $S_{a}$. Suppose instead that there exists $s,s^{\prime}\in\hat{S}_{2}\cap S_{a}$, with $||s-s^{\prime}||\leq\varepsilon^{1/n}$. WLOG assume $s$ was queried after $s^{\prime}$ and let $t$ be the time step on which $s$ was queried. Then $(s^{\prime},\pi^{m}(s^{\prime}))\in X_{t}$. Also, $s,s^{\prime}\in\hat{S}_{2}\cap S_{a}$ implies that and $\pi_{t}(s_{t})=\pi^{m}(s_{t})=a=\pi^{m}(s^{\prime})$. Therefore
$\min_{(s^{\prime\prime},a^{\prime\prime})\in X_{t}:a^{\prime\prime}=\pi_{t}(s_{t})}||s_{t}-s^{\prime\prime}||\leq||s_{t}-s^{\prime}||\leq\varepsilon^{1/n}$
which contradicts $s_{t}\in\hat{S}$. Thus $\hat{S}_{2}\cap S_{a}$ is a $\varepsilon^{1/n}$-packing of $S_{a}$.
By Lemma B.6, there exists a ball $B_{1}$ of diameter $\operatorname*{diam}(\mathbf{s})\sqrt{\frac{n}{2(n+1)}}$ which contains $\mathbf{s}$. Let $R=\operatorname*{diam}(\mathbf{s})\sqrt{\frac{n}{8(n+1)}}$ denote the radius of $B_{1}$. Let $B_{2}$ be the ball with the same center as $B_{1}$ but with radius $\max(R,\varepsilon^{1/n})$. Since $S_{a}\subset\mathbf{s}\subset B_{1}\subset B_{2}$, $\hat{S}_{2}\cap S_{a}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$. Also, $B_{2}$ must contain a ball of radius $\varepsilon^{1/n}$, so Lemma B.5 implies that
$|\hat{S}_{2}\cap S_{a}|\leq$ $\ \mathcal{M}(B_{2},\varepsilon^{1/n})$
@@
$=$ $\ \max\left(\operatorname*{diam}(\mathbf{s})^{n}\left(\frac{n}{8(n+1)}\right)^{n/2},\>\varepsilon\right)\frac{3^{n}}{\varepsilon}$
$\leq$ $\ O\left(\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}+1\right)$
(The $+1$ is necessary for now since $\operatorname*{diam}(\mathbf{s})$ could theoretically be zero.) Therefore
$\operatorname*{\mathbb{E}}[|Q_{T}|]=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{S}|]$
$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{S}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{a\in A}|\hat{S}_{2}\cap S_{a}|\right]$
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+\sum_{a\in A}O\left(\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}+1\right)$
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+|A|\cdot O\left(\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}+1\right)$
$\leq$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\mathbf{s})^{n}}{\varepsilon}\right)$
@@
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{T^{1+\frac{2n}{2n+1}}}+\frac{d}{\sigma}T^{1-\frac{-2n}{2n+1}}\log(T^{\frac{2n}{2n+1}})\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\mathbf{s})^{n}\right)$
$=$ $\ O\left(T^{\frac{2n+0.5}{2n+1}}+\frac{d}{\sigma}T^{\frac{1}{2n+1}}\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\mathbf{s})^{n}\right)$
$\leq$ $\ O\left(T^{\frac{4n+1}{4n+2}}\Big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\Big{)}\right)$
See 5.2.1
Proof.
Let $\xi$ be the event that Algorithm 1 exceeds $T^{\frac{4n+1.5}{4n+2}}(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n})$ queries. Noting that $|Q_{T}|\leq T$ always, we have
$\operatorname*{\mathbb{E}}\big{[}|Q_{T}|\big{]}=$ $\ \Pr[\xi]\,\operatorname*{\mathbb{E}}\big{[}|Q_{T}|\mid\xi\big{]}+\Pr[\neg\xi]\,\operatorname*{\mathbb{E}}\big{[}|Q_{T}|\mid\neg\xi\big{]}$
$\geq$ $\ \Pr[\xi]\,T^{\frac{4n+1.5}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)$
Therefore by Theorem 5.2,
$\Pr[\xi]\leq\frac{O\left(T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)\right)}{T^{\frac{4n+1.5}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)}\leq O\left(T^{\frac{-0.5}{4n+2}}\right)$
If $\xi$ does not occur, then the regret bound from Theorem 5.2 holds. If not, we still have $R_{T}\leq 1$ unconditionally. Thus the regret of the modified version of Algorithm 1 satisfies
$\operatorname*{\mathbb{E}}[R_{T}]\leq$ $\ \Pr[\xi]\,\operatorname*{\mathbb{E}}[R_{T}\mid\xi]+\Pr[\neg\xi]\,\operatorname*{\mathbb{E}}[R_{T}\mid\neg\xi]$
$\leq$ $\ O\left(T^{\frac{-0.5}{4n+2}}\right)\cdot 1+1\cdot O\left(\frac{dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
$\leq$ $\ O\left(\frac{dL}{\sigma}T^{\frac{-1}{8n+4}}\log T\right)$
@@ B.3 Adaptive adversaries
In the standard online learning model, Hedge guarantees sublinear regret against both oblivious and adaptive adversaries (Chapter 5 of [35] or Chapter 21 of [34]). However, [33] state their result only for oblivious adversaries. In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in [33]) must also hold for adaptive adversaries. In this section, we argue why the proof of Lemma 5.4 (Lemma 3.5 in their paper) goes through for adaptive adversaries as well. For this rest of Section B.3, lemma numbers refer to the numbering in [33].
@@
Recall from Section B.1 that [33] allow two separate parameters $k$ and $\hat{k}$, which we unify for simplicity. Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2. More precisely, on each time step $t$, the algorithm samples a Bernoulli random variable $X_{t}\sim\text{Ber}(p)$ and queries if $X_{t}=1$. The key idea is that $X_{t}$ is independent of events on previous time steps. Thus even conditioning on the history up to time $t$, for any for any random variable $Y_{t}$ we can write
@@
This insight immediately extends Observation 3.3 to adaptive adversaries (with the minor modification that queries are now issued independently with probability $p$ on each time step instead of issuing $k$ uniformly distributed queries). Specifically, using the notation from [33] where $i_{t}$ is the action chosen at time $t$, $i_{t}^{0}$ is the action chosen at time $t$ if a query is not issued, and $i_{t}^{*}$ is the optimal action at time $t$, we have
@@
We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of [34]). For each action $a$, we will learn a binary classifier which predicts whether action $a$ is the mentor’s action. Formally, for each $a\in A$, define the policy class $\Pi_{a}=\{\pi_{a}:\pi\in\Pi\text{ and }\pi_{a}(s)=\mathbf{1}(\pi(s)=a))\ \forall s\in S\}$. Informally, for each policy $\pi:S\to A$ in $\Pi$, there exists a policy $\pi_{a}:S\to\{0,1\}$ in $\Pi_{a}$ such that $\pi_{a}(s)=\mathbf{1}(\pi(s)=a)$ for all $s\in S$.
Algorithm 3 runs one copy of our binary-action algorithm Algorithm 1 for each action $a\in A$. At each time step $t$, the copy for action $a$ returns an action $b_{t}^{a}$, with $b_{t}^{a}=1$ indicating a belief that $a=\pi^{m}(s_{t})$ and $b_{t}^{a}=0$ indicating a belief that $a\neq\pi^{m}(s_{t})$. (Note that $b_{t}^{a}=\hat{a}$ is also possible, indicating that the mentor was queried.)
The key idea is that if $b_{t}^{a}$ is correct for each action $a$, there will be exactly one $a$ such that $b_{t}^{a}=1$, and specifically it will be $a=\pi^{m}(s_{t})$. Thus we are guaranteed to take the mentor’s action on such time steps. The analysis for Theorem 5.2 (specifically, Lemma B.1) bounds the number of time steps when a given copy of Algorithm 1 is incorrect, so by the union bound, the number of time steps where any copy is incorrect is $|A|$ times that bound. That in turn bounds the number of time steps where Algorithm 3 takes an action other than the mentor’s. Similarly, the number of queries made by Algorithm 3 is at most $|A|$ times the bound from Theorem 5.2. The result is the following theorem:
@@
4:              $\tilde{\Pi}_{a}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
@@
6:              $\tilde{\Pi}_{a}\leftarrow$ any adversarial $\varepsilon$-cover of size at most $(eT/d)^{d}$
@@
8:         for $a\in A$ do
9:              $b_{t}^{a}\leftarrow$ action from running one step of Algorithm 1 on $\Pi_{a}$ (with the same $T,\varepsilon,d$)
10:         if $b_{t}^{a}\neq\hat{a}$ $\forall a\in A$ and $\exists a\in A:b_{t}^{a}=1$ then
11:              Take any action $a$ with $b_{t}^{a}=1$
@@
Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{a}$ has finite VC dimension $d$ and $\mathbf{s}$ is $\sigma$-smooth or (2) $\Pi_{a}$ has finite Littlestone dimension $d$ for all $a\in A$. Then for any $T\in\mathbb{N}$, Algorithm 3 with $T$ and $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{|A|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(|A|T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\right)\right)$
@@
    We refer to the copy of Algorithm 1 running on $\Pi_{a}$ as “copy $a$ of Algorithm 1”.
@@
    Let $\pi_{t}^{a}$ and $X_{t}^{a}$ refer to the values of $\pi_{t}$ and $X_{t}$ for copy $a$ of Algorithm 1.
@@
    Let $\pi^{ma}:S\to\{0,1\}$ be the policy defined by $\pi^{ma}(s)=\mathbf{1}(\pi^{m}(s_{t})=a)$. Note that querying the mentor tells the agent $\pi^{m}(s_{t})$, which allows the agent to compute $\pi^{ma}(s_{t})$: this is necessary when Algorithm 1 queries while running on some $\Pi_{a}$.
@@
    Let $V_{T}^{a}=\{t\in[T]:b_{t}^{a}\neq\pi^{ma}(s_{t})\}$ be the set of time steps where $\pi_{t}^{a}$ does not correctly determine whether the mentor would take action $a$ and let $V_{T}=\{t\in[T]:a_{t}\neq\pi^{m}(s_{t})\}$ be the set of time steps where the agent’s action doesn’t match the mentor’s.
@@
We claim that $V_{T}\subseteq\cup_{a\in A}V_{T}^{a}$. Suppose the opposite: then there exists $t\in V_{T}$ such that $b_{t}^{a}=\pi^{ma}(s_{t})$ for all $a\in A$. Since $\pi^{m}(s_{t})\in A$, there is exactly one $a\in A$ such that $\mathbf{1}(\pi^{m}(s_{t})=a)=\pi^{ma}(s_{t})=b_{t}^{a}=1$. Specifically, this holds for $a=\pi^{m}(s_{t})$. But then Algorithm 3 takes action $\min\{a\in A:b_{t}^{a}=1\}=\pi^{m}(s_{t})$, which contradicts $t\in V_{T}$. Therefore $V_{T}\subseteq\cup_{a\in A}V_{T}^{a}$, and applying the union bound completes the proof. ∎
@@
The argument is similar to the proof of Lemma B.2. If $\mu^{m}(s_{t})\neq\mu(s_{t},a_{t})$, then $a_{t}=a$ for some $a\in A$ where $b_{t}^{a}=1$. Therefore copy $a$ of Algorithm 1 did not query at time $t$ and $\pi_{t}^{a}(s_{t})=1$. Let $(s^{\prime},a^{\prime})=\operatorname*{arg\,min}_{(s,a)\in X_{t}^{a}:\pi_{t}^{a}(s_{t})=a}||s_{t}-s||$. Then $||s_{t}-s^{\prime}||\leq\varepsilon^{1/n}$ and $a^{\prime}=\pi_{t}^{a}(s_{t})=1$.
By construction of $X_{t}^{a}$, $a^{\prime}=\pi^{ma}(s^{\prime})$ so $\pi^{ma}(s^{\prime})=1$ which implies $\pi^{m}(s^{\prime})=a$. Then by the local generalization assumption,
$\mu(s_{t},a_{t})=\mu(s_{t},a)=\mu(s_{t},\pi^{m}(s^{\prime}))\geq\mu^{m}(s_{t})-L||s_{t}-s^{\prime}||\geq\mu^{m}(s_{t})-L\varepsilon^{1/n}$
@@
Theorem 5.2 implies that each copy of Algorithm 1 makes $O\big{(}T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\big{)}\big{)}$ queries in expectation, so by linearity of expectation, the expected number of queries made by Algorithm 3 is $O\big{(}|A|T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\mathbf{s})^{n}\big{)}\big{)}$. Using the same argument as in the proof of Lemma B.7 (with Lemma C.2 replacing Lemma B.2), we get
$R_{T}=$ $\ \prod_{t=1}^{T}\mu^{m}(s_{t})-\prod_{t=1}^{T}\mu(s_{t},a_{t})$
$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu^{m}(s_{t})-\min(\mu^{m}(s_{t}),\mu(s_{t},a_{t}))\Big{)}$
$\leq$ $\ |V_{T}|L\varepsilon^{1/n}$
Then by Lemma C.1, $R_{T}\leq L\varepsilon^{1/n}\sum_{a\in A}|V_{T}^{a}|$. Taking the expectation and applying Lemma B.1 to each $V_{T}^{a}$ gives us
$\operatorname*{\mathbb{E}}[R_{T}]\leq L\varepsilon^{1/n}\sum_{a\in A}O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)=O\left(|A|L\varepsilon^{1/n}\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
as required. ∎

### v4

@@ B.2 Main proof
Next, fix an $y\in\mathcal{Y}$ and let $X_{y}=\{x\in\boldsymbol{x}:\pi^{m}(x)=y\}$ be the set of observed inputs which share a mentor action. We claim that $\hat{X}_{2}\cap X_{y}$ is a packing of $X_{y}$. Suppose instead that there exists $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$, with $||x-x^{\prime}||\leq\varepsilon^{1/n}$. WLOG assume $x$ was queried after $x^{\prime}$ and let $t$ be the time step on which $x$ was queried. Then $(x^{\prime},\pi^{m}(x^{\prime}))\in X_{t}$. Also, $x,x^{\prime}\in\hat{X}_{2}\cap X_{y}$ implies that and $\pi_{t}(x_{t})=\pi^{m}(x_{t})=y=\pi^{m}(x^{\prime})$. Therefore
$\min_{(x^{\prime\prime},y^{\prime\prime})\in X_{t}:y^{\prime\prime}=\pi_{t}(x_{t})}||x_{t}-x^{\prime\prime}||\leq||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$
which contradicts $x_{t}\in\hat{X}$. Thus $\hat{X}_{2}\cap X_{y}$ is a $\varepsilon^{1/n}$-packing of $X_{y}$.
By Lemma B.6, there exists a ball $B_{1}$ of diameter $\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{2(n+1)}}$ which contains $\boldsymbol{x}$. Let $R=\operatorname*{diam}(\boldsymbol{x})\sqrt{\frac{n}{8(n+1)}}$ denote the radius of $B_{1}$. Let $B_{2}$ be the ball with the same center as $B_{1}$ but with radius $\max(R,\varepsilon^{1/n})$. Since $X_{y}\subset\boldsymbol{x}\subset B_{1}\subset B_{2}$, $\hat{X}_{2}\cap X_{y}$ is also a $\varepsilon^{1/n}$-packing of $B_{2}$. Also, $B_{2}$ must contain a ball of radius $\varepsilon^{1/n}$, so Lemma B.5 implies that
$|\hat{X}_{2}\cap X_{y}|\leq$ $\ \mathcal{M}(B_{2},\varepsilon^{1/n})$
@@
$=$ $\ \max\left(\operatorname*{diam}(\boldsymbol{x})^{n}\left(\frac{n}{8(n+1)}\right)^{n/2},\>\varepsilon\right)\frac{3^{n}}{\varepsilon}$
$\leq$ $\ O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
(The $+1$ is necessary for now since $\operatorname*{diam}(\boldsymbol{x})$ could theoretically be zero.) Therefore
$\operatorname*{\mathbb{E}}[|Q_{T}|]=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{X}|]$
$=$ $\ \sqrt{\frac{T}{\varepsilon}}+\operatorname*{\mathbb{E}}[|\hat{X}_{1}|]+\operatorname*{\mathbb{E}}\left[\sum_{y\in\mathcal{Y}}|\hat{X}_{2}\cap X_{y}|\right]$
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+\sum_{y\in\mathcal{Y}}O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
$\leq$ $\ \sqrt{\frac{T}{\varepsilon}}+O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)+|\mathcal{Y}|\cdot O\left(\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}+1\right)$
$\leq$ $\ O\left(\sqrt{\frac{T}{\varepsilon}}+\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T+\frac{\operatorname*{diam}(\boldsymbol{x})^{n}}{\varepsilon}\right)$
@@
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(\sqrt{T^{1+\frac{2n}{2n+1}}}+\frac{d}{\sigma}T^{1-\frac{-2n}{2n+1}}\log(T^{\frac{2n}{2n+1}})\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\boldsymbol{x})^{n}\right)$
$=$ $\ O\left(T^{\frac{2n+0.5}{2n+1}}+\frac{d}{\sigma}T^{\frac{1}{2n+1}}\log T+T^{\frac{2n}{2n+1}}\operatorname*{diam}(\boldsymbol{x})^{n}\right)$
$\leq$ $\ O\left(T^{\frac{4n+1}{4n+2}}\Big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\Big{)}\right)$
If we instead use the second bound from Lemma B.4, the same arithmetic gives us:
See 5.3
@@ B.3 Adaptive adversaries
In the standard online learning model, Hedge guarantees sublinear regret against both oblivious and adaptive adversaries (Chapter 5 of Slivkins et al. (2019) or Chapter 21 of Shalev-Shwartz & Ben-David (2014)). However, Russo et al. (2024) state their result only for oblivious adversaries. In order for our overall proof of Theorem 5.1 to hold for adaptive adversaries, Lemma 5.4 (Lemma 3.5 in Russo et al. (2024)) must also hold for adaptive adversaries. In this section, we argue why the proof of Lemma 5.4 (Lemma 3.5 in their paper) goes through for adaptive adversaries as well. For this rest of Section B.3, lemma numbers refer to the numbering in Russo et al. (2024).
@@
Recall from Section B.1 that Russo et al. (2024) allow two separate parameters $k$ and $\hat{k}$, which we unify for simplicity. Recall also that Lemma 3.5 refers to the variant of Hedge which queries with probability $p=\hat{k}/T=k/T$ independently on each time step (Algorithm 2. More precisely, on each time step $t$, the algorithm samples a Bernoulli random variable $X_{t}\sim\text{Ber}(p)$ and queries if $X_{t}=1$. The key idea is that $X_{t}$ is independent of events on previous time steps. Thus even conditioning on the history up to time $t$, for any for any random variable $Y_{t}$ we can write
@@
This insight immediately extends Observation 3.3 to adaptive adversaries (with the minor modification that queries are now issued independently with probability $p$ on each time step instead of issuing $k$ uniformly distributed queries). Specifically, using the notation from Russo et al. (2024) where $i_{t}$ is the action chosen at time $t$, $i_{t}^{0}$ is the action chosen at time $t$ if a query is not issued, and $i_{t}^{*}$ is the optimal action at time $t$, we have
@@
We use the standard “one versus rest” reduction (see, e.g., Chapter 29 of Shalev-Shwartz & Ben-David (2014)). For each action $y$, we will learn a binary classifier which predicts whether action $y$ is the mentor’s action. Formally, for each $y\in\mathcal{Y}$, define the policy class $\Pi_{y}=\{\pi_{y}:\pi\in\Pi\text{ and }\pi_{y}(x)=\mathbf{1}(\pi(x)=y))\ \forall x\in\mathcal{X}\}$. In words, for each policy $\pi:\mathcal{X}\to\mathcal{Y}$ in $\Pi$, there exists a policy $\pi_{y}:\mathcal{X}\to\{0,1\}$ in $\Pi_{y}$ such that $\pi_{y}(x)=\mathbf{1}(\pi(x)=y)$ for all $x\in\mathcal{X}$.
Algorithm 3 runs one copy of our binary-action algorithm Algorithm 1 for each action $y\in\mathcal{Y}$. At each time step $t$, the copy for action $y$ returns an action $b_{t}^{y}$, with $b_{t}^{y}=1$ indicating a belief that $y=\pi^{m}(x_{t})$ and $b_{t}^{y}=0$ indicating a belief that $y\neq\pi^{m}(x_{t})$. (Note that $b_{t}^{y}=\tilde{y}$ is also possible, indicating that the mentor was queried.)
The key idea is that if $b_{t}^{y}$ is correct for each action $y$, there will be exactly one $y$ such that $b_{t}^{y}=1$, and specifically it will be $y=\pi^{m}(x_{t})$. Thus we are guaranteed to take the mentor’s action on such time steps. The analysis for Theorem 5.2 (specifically, Lemma B.1) bounds the number of time steps when a given copy of Algorithm 1 is incorrect, so by the union bound, the number of time steps where any copy is incorrect is $|\mathcal{Y}|$ times that bound. That in turn bounds the number of time steps where Algorithm 3 takes an action other than the mentor’s. Similarly, the number of queries made by Algorithm 3 is at most $|\mathcal{Y}|$ times the bound from Theorem 5.2. The result is the following theorem:
@@
4:              $\tilde{\Pi}_{y}\leftarrow$ any smooth $\varepsilon$-cover of $\Pi$ of size at most $(41/\varepsilon)^{d}$
@@
6:              $\tilde{\Pi}_{y}\leftarrow$ any adversarial $\varepsilon$-cover of size at most $(eT/d)^{d}$
@@
8:         for $y\in\mathcal{Y}$ do
9:              $b_{t}^{y}\leftarrow$ action from running one step of Algorithm 1 on $\Pi_{y}$ (with the same $T,\varepsilon,d$)
10:         if $b_{t}^{y}\neq\tilde{y}$ $\forall y\in\mathcal{Y}$ and $\exists a\in\mathcal{Y}:b_{t}^{y}=1$ then
11:              Take any action $y$ with $b_{t}^{y}=1$
@@
Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{y}$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi_{y}$ has finite Littlestone dimension $d$ for all $y\in\mathcal{Y}$. Then for any $T\in\mathbb{N}$, Algorithm 3 with $T$ and $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
$\operatorname*{\mathbb{E}}\left[R_{T}\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$
$\operatorname*{\mathbb{E}}[|Q_{T}|]\in$ $\ O\left(|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\left(\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\right)\right)$
@@
    We refer to the copy of Algorithm 1 running on $\Pi_{y}$ as “copy $y$ of Algorithm 1”.
@@
    Let $\pi_{t}^{y}$ and $X_{t}^{y}$ refer to the values of $\pi_{t}$ and $X_{t}$ for copy $y$ of Algorithm 1.
@@
    Let $\pi^{my}:\mathcal{X}\to\{0,1\}$ be the policy defined by $\pi^{my}(x)=\mathbf{1}(\pi^{m}(x_{t})=y)$. Note that querying the mentor tells the agent $\pi^{m}(x_{t})$, which allows the agent to compute $\pi^{my}(x_{t})$: this is necessary when Algorithm 1 queries while running on some $\Pi_{y}$.
@@
    Let $M_{T}^{y}=\{t\in[T]:b_{t}^{y}\neq\pi^{my}(x_{t})\}$ be the set of time steps where $\pi_{t}^{y}$ does not correctly determine whether the mentor would take action $y$ and let $M_{T}=\{t\in[T]:y_{t}\neq\pi^{m}(x_{t})\}$ be the set of time steps where the agent’s action doesn’t match the mentor’s.
@@
We claim that $M_{T}\subseteq\cup_{y\in\mathcal{Y}}M_{T}^{y}$. Suppose the opposite: then there exists $t\in M_{T}$ such that $b_{t}^{y}=\pi^{my}(x_{t})$ for all $y\in\mathcal{Y}$. Since $\pi^{m}(x_{t})\in\mathcal{Y}$, there is exactly one $y\in\mathcal{Y}$ such that $\mathbf{1}(\pi^{m}(x_{t})=y)=\pi^{my}(x_{t})=b_{t}^{y}=1$. Specifically, this holds for $y=\pi^{m}(x_{t})$. But then Algorithm 3 takes action $\min\{y\in\mathcal{Y}:b_{t}^{y}=1\}=\pi^{m}(x_{t})$, which contradicts $t\in M_{T}$. Therefore $M_{T}\subseteq\cup_{y\in\mathcal{Y}}M_{T}^{y}$, and applying the union bound completes the proof. ∎
@@
The argument is similar to the proof of Lemma B.2. If $\mu_{t}^{m}(x_{t})\neq\mu_{t}(x_{t},y_{t})$, then $y_{t}=y$ for some $y\in\mathcal{Y}$ where $b_{t}^{y}=1$. Therefore copy $y$ of Algorithm 1 did not query at time $t$ and $\pi_{t}^{y}(x_{t})=1$. Let $(x^{\prime},y^{\prime})=\operatorname*{arg\,min}_{(x,y)\in X_{t}^{y}:\pi_{t}^{y}(x_{t})=y}||x_{t}-x||$. Then $||x_{t}-x^{\prime}||\leq\varepsilon^{1/n}$ and $y^{\prime}=\pi_{t}^{y}(x_{t})=1$.
By construction of $X_{t}^{y}$, $y^{\prime}=\pi^{my}(x^{\prime})$ so $\pi^{my}(x^{\prime})=1$ which implies $\pi^{m}(x^{\prime})=y$. Then by the local generalization assumption,
$\mu_{t}(x_{t},y_{t})=\mu_{t}(x_{t},y)=\mu_{t}(x_{t},\pi^{m}(x^{\prime}))\geq\mu_{t}^{m}(x_{t})-L||x_{t}-x^{\prime}||\geq\mu_{t}^{m}(x_{t})-L\varepsilon^{1/n}$
@@
Theorem 5.2 implies that each copy of Algorithm 1 makes $O\big{(}T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\big{)}\big{)}$ queries in expectation, so by linearity of expectation, the expected number of queries made by Algorithm 3 is $O\big{(}|\mathcal{Y}|T^{\frac{4n+1}{4n+2}}\big{(}\frac{d}{\sigma}\log T+\operatorname*{diam}(\boldsymbol{x})^{n}\big{)}\big{)}$. Using the same argument as in the proof of Lemma B.4 (with Lemma C.2 replacing Lemma B.2), we get
$R_{T}=$ $\ \prod_{t=1}^{T}\mu_{t}^{m}(x_{t})-\prod_{t=1}^{T}\mu_{t}(x_{t},y_{t})$
$\leq$ $\ \sum_{t=1}^{T}\Big{(}\mu_{t}^{m}(x_{t})-\min(\mu_{t}^{m}(x_{t}),\mu_{t}(x_{t},y_{t}))\Big{)}$
$\leq$ $\ |M_{T}|L\varepsilon^{1/n}$
Then by Lemma C.1, $R_{T}\leq L\varepsilon^{1/n}\sum_{y\in\mathcal{Y}}|M_{T}^{y}|$. Taking the expectation and applying Lemma B.1 to each $M_{T}^{y}$ gives us
$\operatorname*{\mathbb{E}}[R_{T}]\leq L\varepsilon^{1/n}\sum_{y\in\mathcal{Y}}O\left(\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)=O\left(|\mathcal{Y}|L\varepsilon^{1/n}\frac{d}{\sigma}T\varepsilon\log(1/\varepsilon)\log T\right)$
Plugging in $\varepsilon=T^{\frac{-2n}{2n+1}}$ produces the desired bounds (see the arithmetic in the proof of Theorem 5.2 in Appendix B). ∎
One can use the same argument to obtain a multi-action additive regret bound: simply modify the proof of Theorem C.1 to use $\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\leq|M_{T}|L\varepsilon^{1/n}$ instead of $R_{T}\leq|M_{T}|L\varepsilon^{1/n}$.
Corollary C.1.1.
Assume $\pi^{m}\in\Pi$ where either (1) $\Pi_{y}$ has finite VC dimension $d$ and $\boldsymbol{x}$ is $\sigma$-smooth or (2) $\Pi_{y}$ has finite Littlestone dimension $d$ for all $y\in\mathcal{Y}$. Then for any $T\in\mathbb{N}$, Algorithm 3 with $T$ and $\varepsilon=T^{\frac{-2n}{2n+1}}$ satisfies
$\operatorname*{\mathbb{E}}\left[\sum_{t=1}^{T}\mu_{t}^{m}(x_{t})-\sum_{t=1}^{T}\mu_{t}(x_{t},y_{t})\right]\in$ $\ O\left(\frac{|\mathcal{Y}|dL}{\sigma}T^{\frac{-1}{2n+1}}\log T\right)$